OpenStudy (anonymous):

Eric measured the height of a window frame as 5.43 feet, but the actual height was 6 feet. What is the percentage of error in Eric's measurement? 9.50% 10.52% 11.28% 12.43%

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Elsa213 @dan815 @BioHazard9064

OpenStudy (anonymous):

Start by setting up a fraction

OpenStudy (anonymous):

Since six is the whole, it will be your base, just like saying the frame was 5.43 feet out of six

OpenStudy (anonymous):

I get that part but would u have to multiply?

OpenStudy (anonymous):

Well let's look at an easier version of the problem. If I have 67 cents, what percent of a dollar do I have?

OpenStudy (anonymous):

My answer was c btw

OpenStudy (anonymous):

That's not what I got

OpenStudy (anonymous):

0% ?

OpenStudy (anonymous):

Well there are 100 cents in one dollar, and we have 67 out of 100, or 67/100

OpenStudy (anonymous):

In decimal form that is 0.67. To convert to a percent we would move the decimal place to the right two points, or multiply by 100

OpenStudy (anonymous):

Your problem is similar, except we're looking for a percent difference between his measurement and the actual dimensions. So we can say 5.43/6

OpenStudy (anonymous):

This will give you what percent of the window he measured, to get the percent difference, you can subtract that answer from 100. You should get 9.5%

OpenStudy (phi):

***5.43 feet, but the actual height was 6 feet. What is the percentage of error*** before finding the percent error, you need to find the error in feet For example, if he measured 6 ft he would be exactly right, and have 0 error if he measured 5 ft, he would be off by 1 foot. the error would be 1 ft here the error is the difference between 6 and 5.43 6.00 - 5.43 = 0.57 now we can find percent error: first find the fraction error/true 0.57/6.00 second, change that to a decimal. (hopefully you know how to do that) you get 0.095 third, change a decimal to a percent. (per cent is Latin for "per one hundred") multiply top and bottom by 100: \[ \frac{0.095}{1} \cdot \frac{100}{100} = \frac{9.5}{100} \] that shows 9.5 per hundred, which we can also write as 9.5% (the % means per hundred, i.e. divide by 100)